OpenStudy (anonymous):

How do you solve this set of linear equations? 40x + 20y = 5000 25x + 50y = 1000

OpenStudy (anonymous):

use elimination

OpenStudy (anonymous):

\[40x+20y=5000\] \[25x+50y=1000\] Now multiply the first equation by -5 and the second equation by 2 \[40x+20y=5000(-5)\] \[25x+50y=1000(2)\]

OpenStudy (anonymous):

\[-200x-100y=-25,000\] \[50x+100y=2000\] the -100y and 100y cancel out

OpenStudy (anonymous):

now bring down the -200x by subtracting 50x. then bring down the -25,000 by subtracting 2000

OpenStudy (anonymous):

Thanks now you have the equation \[-250x=-27,000\] solve for x to get 108

OpenStudy (anonymous):

now plug 108 back into the original equation 40x+20y=5000 solve for y: \[40(108)+20y=5000\] \[4320+20y=5000\] Subtract 4320 from both sides \[20y=680\] y=20

OpenStudy (anonymous):

Do you understand @kennybm

OpenStudy (anonymous):

Yeah I'm getting it, hang on :)

OpenStudy (anonymous):

ok ;)

OpenStudy (anonymous):

was it correct?